OpenStudy (anonymous):

2over w=xover 5;w solve the formula plz help

OpenStudy (anonymous):

\[\frac{ 2 }{ w } = \frac{ x }{ 5 }\] is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cross multiply. w=10/x

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):